OpenStudy (anonymous):

The height of an arch is given by the formula h(x) = - 3/64 x raised to the power of 2 +27, where x is the horizontal distance from the center of the arch? a) What is the maximum height of the arch? b) What is the height of the arch 10ft from the center? c) How far from the center is the arch 8ft high?

OpenStudy (dumbcow):

to maximize h, differentiate and set equal to 0

OpenStudy (dumbcow):

for part b) substitute 10 for x for part c) substitute 8 for h and solve for x

OpenStudy (anonymous):

what is the answers when you differentiate? I am unsure

OpenStudy (dumbcow):

given f(x) = x^n f'(x) = n*x^(n-1) apply this to your function -3/64 x^2 +27

OpenStudy (anonymous):

this is confusing can't you work it out I don't understand. Thanks